OpenStudy (anonymous):

Find the value that completed the square and then rewrite as a perfect square. z^2-19z+_

OpenStudy (agent0smith):

To complete the square for ax^2 + bx + _, just divide b by 2, and square that. So here, first divide -19 by 2, then square that result.

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

but wait you can't square root a negitive number?

OpenStudy (agent0smith):

You're not square rooting it. You're squaring it.

OpenStudy (anonymous):

oh okay, my bad

OpenStudy (anonymous):

The answer was supposed to be 361/4;(z-19/2)^2 .. and I didn't get that when I tried..

OpenStudy (agent0smith):

That's because that's the final answer. The first step is just getting this:\[\Large \left( \frac{ -19 }{ 2 } \right)^2\]

OpenStudy (anonymous):

how did they get the 361/4?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Square this and find out: \[\Large \left( \frac{ -19 }{ 2 } \right)^2 = \frac{ (-19)^2 }{ 2^2 }\]